how do u do graphs for the follwing equation:
x-y=4
x+y=2

Answer by thechamp1011(19)

To solve this problem, you need to solve x and y interms of each other by the method of substitution:
since, x-y=4, we can rewrite this equation as : x=4+y
then solve the second equattion : x+Y=2
but now you can plug in the new equation for x, then you will get : 4+y+y=2
this becomes: 2y=2-4
2y=-2
therefore, y=-2/2=-1
now we can go back to the original equation: x-y=4
and then plug in the value for the y: x-(-1)=4
the x becomes: x=4-1=5
After you get the both values of x and y, then draw the x and y axis and plot the graph using the x and y point and then connet it together.